OpenStudy (anonymous):

Homework Help (7th grade) thanks

OpenStudy (anonymous):

please explain how you did it

OpenStudy (anonymous):

help pleasee :c

OpenStudy (anonymous):

-20-15 = -35 5-(-2)=7 4-(-10)=14 -40-60=100

OpenStudy (anonymous):

ooo thanks :)))

OpenStudy (anonymous):

can you explain how u did that please?

OpenStudy (anonymous):

is it clear how that works ?

OpenStudy (anonymous):

which example do you want to go over ?

OpenStudy (anonymous):

just do the first one thanks

OpenStudy (anonymous):

we have the numbers -20 and 15. we are asked to add/subtract them, and end up with -35 . If we start with -20 , and subtract 15 more from that, we will end with -35. So: -20 - 15 = -35

OpenStudy (anonymous):

ooh ok so basically we are just finding out if we need to add/subtract them and then figure it out?

OpenStudy (anonymous):

Looks to me like that's what you're supposed to do

OpenStudy (anonymous):

ok thanks :))

OpenStudy (anonymous):

y'absolutely :)